Search

Date Entry File Actions
2025-02-27

Replaced weather station battery at Hancock. Tagged instrumetations and some sensors.

Almost no snow.

2025-02-26

Replaced the weather station batteris at Escanaba and Rhinelander. Tagged instruments and some sensors with barcodes.

Minimal snow at either location.

2025-02-25

Replaced the Battery at the Lake City MLE weather station and logger 125. Labeled sensors with qrcodes.

Snow depth was 15 inches at the snow depth sensor.

2024-12-09

Replaced Battery at the Lake City Weather station. Cleaned out raingauge. 1 extra tip. Checked moisture logger. The 121 logger box lock was jammed. Need to bring some graphite next time.

2024-11-20

Lux marginal land site visited by Kevin/Stacey today. Found many corner posts and sockets pulled out and also north side of G10R3 found soil moisture probes dug up and chewed. Will not be able to repair. They are soil moisture profiler sensors for this plot..

2024-09-24

MLE plant species composition, Lux Arbor (MS) - Alyssa, Slater, Emily, Brian, Olivia, Trajana

2024-09-06

MLE plant species composition, Lake City (MC) - Alyssa, Slater, Emily, Eric

2024-09-05

MLE plant species composition, Escanaba (MN) - Alyssa, Slater, Emily, Eric

2024-09-04

MLE plant species composition, Rhinelander (WN) - Alyssa, Slater, Emily, Eric

2024-09-03

MLE plant species composition, Hancock (WC) - Alyssa, Slater, Emily, Eric

2024-07-01

I replaced the Solar Radiation sensor at Rhinelander with a new Apogee SP230-0SS serial number 12575.

I put a replacement radio on the G5 logger closest to the road.

2024-04-04

Disassembled Lake City shelters on 4/1/24, and Escanaba AND Rhinelander shelters on 4/2/24. Todd decided NOT to bring the trailer to haul back the shelter pieces but will pick them up in the next month. Did not take down Hancock shelters due to an unexpected snow storm. Recalled G5R3 logger was not locked and did so on 4/4. Not sure when it was unlocked, but the logger box door was also unlocked, units seemed fine. Neither Kevin or Sven recall leaving this unlocked. Nothing missing. As Sven mentioned on 4/2, we sampled soil at a few plots in Rhinelander for Ashim Datta.

2024-04-02

RF407 radio on G10 116 was not working at escanaba. I moved it to 119 and it did not work there either. Took the radio back for repair.

Replaced Battery on the weather station at Escanaba

2024-04-02

KK, DW, SB, TM, CR disassembled the four shelters at at Escanaba and four shelters at Rhinelander. Replaced weather station battery at Escanaba and Rhinelander. Took samples for Ashim at Rhinelander.

2024-04-01

KK, DW, SB, TM, CR disassembled the four shelters at Lake City. Stack two shelter sets on one pallet. Todd and Chris will pick up later and return. We removed R1G5 10 and 25cm sensors in the under shelter and ambient plot. Left the remainder for removal another time.

2024-04-01

Replaced Weather station battery at Lake City.

2023-10-11

MLE MC surface soil (0-25cm) sampling - Todd, Chris

2023-10-09

MLE MN surface soil (0-25cm) sampling - Todd, Chris

2023-10-08

Weigh MLE WC soil (10/2/2023) for extraction and moisture - Sidney, Alex

2023-10-03

MLE WN surface soil (0-25cm) sampling - Todd, Chris

2023-10-02

MLE WC surface soil (0-25cm) sampling - Todd, Chris

2023-09-20

MLE species composition in Lux Arbor, Michigan South (MS). Two transects through the fertilized portion and 2 in the unfertilized. Stopping every ~1m to identify and measure the height of the first intercepted plant. - Alyssa, Evan, Gloria, Helen, Rick

2023-09-18

Removed flags and chambers from Martha’s study on the north baitcrop area BCSE.

Additionally, removed all shelters on the Lux Arbor marginal land site. Three reps show high stress. Also, deer pellets in three of the four sheltered areas. ALOT! Only deployed one set of RH/T sensors in rep four this year.

2023-09-15

Lake City Bulk Density cores at all plots 0-4” and 5-9” depths. Shelters moved into winter positions and screws deployed. Checked all loggers G5r2 121 SN 43047, G5r1 127 SN43040, weather SN37075, G5r3 125 SN43043. Checked rain gauge. Some labels were missing, written on core or cap. Possibly sampled a few on an earlier trip with Ashim. Removed RH/T sensors around noon.

2023-09-14

MLE species composition in Lake City, Michigan Central (MC). Two transects through the fertilized portion and 2 in the unfertilized. Stopping every ~1m to identify and measure the height of the first intercepted plant. - Alyssa, Evan, Gloria, Helen

2023-09-14

Sampled bulk density at all Escanaba plots 0-4” and 5-9” depths, 2” wide slide hammer sampler AMS, Inc. Checked all loggers G5r1 116 SN42921, G10r3 119 SN43039, G5r2 120 SN43044, 38603 weather station. Removed all shelters by noon. Removed RH/T sensors. Checked rain gage

2023-09-13

Sampled Rhinelander all plots for Bulk Density. 0-4” and 5-9” depth cores (2” wide). Also sampled bulk soil at 22 plots, both depths. Soil was rocky so we used a drain spade shove for this portion. Filled in all holes. Checked data loggers and recorded SNs. SN 38838 weather station, G5r1 126 SN43048, G5r3 124 SN43041, G5r2 118 SN43042. Checked rain gage. Filled in all holes.

2023-09-13

MLE species composition in Escanaba, Michigan North (MN). Two transects through the fertilized portion and 2 in the unfertilized. Stopping every ~1m to identify and measure the height of the first intercepted plant. - Alyssa, Evan, Gloria, Helen

2023-09-12

Sampled all plots for bulk density. 0-4” and 5-9”. One inch spacing is due to the non-sampled portion of first core. That one inch is discarded. Checked all loggers and recorded SNs. G5r4 128 SN43046, G5r3 129 SN43045, G5r2 122 SN43049, SN 38839 weather station.

2023-09-12

MLE species composition in Rhinelander, Wisconsin Central (WN). Two transects through the fertilized portion and 2 in the unfertilized. Stopping every ~1m to identify and measure the height of the first intercepted plant. - Alyssa, Evan, Gloria, Helen

2023-09-11

MLE species composition in Hancock, Wisconsin Central (WC). Two transects through the fertilized portion and 2 in the unfertilized. Stopping every ~1m to identify and measure the height of the first intercepted plant. - Alyssa, Evan, Gloria, Helen

2023-09-06

Lake City Shelters removed by Kravchenko lab R1 3:00PM R2 2:45PM R3 3:25PM R4 3:40PM

2023-08-30

Set fallen HOBO sensor on t-post at Lux Arbor MLE switchgrass site (block 3?). Sensor remained about 1m off ground while collecting previous data (unsure when sensor fell over sideways). Now is upright, approximately 20-30 cm above switchgrass. Inside shelter sensor remains solid.

2023-07-20
  • 7/17 - Lake City with Dave and Ashim to collect bulk soil samples at 0-10 and 10-25cm using a 2” sampler and slide hammer (at all sites). Bulk density will be collected in early fall.
  • 7/18 - Escanaba and
  • 7/19 Rhinelander and Hancock. Rhinelander soil very rocky and dry, will need to finish R3 and 4 Bulk soil and bulk density.

Changed battery at weather station Rhinelander. Changed PAR and Solar Radiation sensor at Lake City and Escanaba. Dave installed RH/T and two shelters each at Lake City and Escanaba., Calibration images saved here.

2023-07-19

Rhinelander and Hancock bulk soil sampling KK, DW, AD. Three 0-4” and 4-10” cores per plot. All holes filled. All plots sampled.

2023-07-19

Sampled all plots for bulk soil at Rhinelander and Hancock MLEs. Three cores each of 0-10cm then 10-25cm below. Each core was taken in a one meter square area, one step north of the plot centerline. No more than two meters in the plot (from side) Cores were collected about .5-1 m apart. Holes filled with soil just outside or at the edge of the plot. Took approximately five hours. KK, DW, AD. 2” core samplers. No bulk density. G5 cores taken 7 steps just outside the reserve area. All collected in +N. New battery at the weather station in Rhinelander. We broke two samplers at Rhinelander and stopped bulk soil just prior to finishing R2. All poplars sampled. Used a drain spade to sample at R3 and 4 in Hancock. Created an approximately 4” diameter circle, then inserted spade 10cm and popped up soil. cleared hole, repeated in the 10-25cm depth. Worked very well and was quite fast. Filled in holes when completed.

2023-07-18

Sampled all plots for bulk soil at Escanaba MLE. Three cores each of 0-10cm then 10-25cm below. Each core was taken in a one meter square area, one step north of the plot centerline. No more than two meters in the plot (from side) Cores were collected about .5-1 m apart. Holes filled with soil just outside or at the edge of the plot. Took approximately five hours. KK, DW, AD. 2” core samplers. No bulk density. G5 cores taken 7 steps just outside the reserve area. All collected in +N. Also added two sets of RH/T sensors (one under shelter and one in ambient 0N on a post). Swapped PAR and solar radiation sensors for recalibration.

2023-07-18

Escanaba bulk soil sampling KK, DW, AD. Three 0-4” and 4-10” cores per plot. All holes filled. All plots sampled.

2023-07-17

Sampled all plots for bulk soil at Lake City MLE. Three cores each of 0-10cm then 10-25cm below. Each core was taken in a one meter square area, one step east of the plot centerline. No more than two meters in the plot (from side) Cores were collected about .5-1 m apart. Holes filled with soil just outside or at the edge of the plot. Took approximately five hours. KK, DW, AD. 2” core samplers. No bulk density. G5 cores taken 7 steps just outside the reserve area. All collected in +N. Also added two sets of RH/T sensors (one under shelter and one in ambient 0N on a post). Swapped PAR and solar radiation sensors for recalibration.

2023-07-17

Lake City bulk soil sampling for C and N. Ashim has SURF. Three cores 0-4” and 4-10”, collected 3 per plot of each depth. Bagged and labeled. KK, DW, AD. Very dry conditions.

2023-05-30

Chris and Todd installed shelters at Escanaba and Lake City after fertilization of G5. Will install Lux Arbor tomorrow (Kevin and Dave).

2023-03-24

Hancock replaced by: 3/22 Apogee SP230-L. WO#275051 5s may be 6s SO# 323936 mult 5 SN7775 or 7776 Licor Quantum Q116919. Licor part # li190r-L. ButQ116920 on cable? Removed for calibration Q101964 quantum Apogee SP230-L mult 5.0 sn7014 New weather station battery.

Rhinelander 3/23 SP230-L sn 7772 mult 5.0 apogee installed SP230-L sn. 6347 Removed Licor Quantum an 103555 LI190R installed Li190R. sn. 102140 Removed Rhinelander G8 camera not located. Too much snow. G5 logger not transmitting. Downloaded data. G5 camera never turned on. New battery weather station.

Escanaba. 3/23 G5 camera and post gone. C10 camera on the ground. New female terminals on G10 logger. New battery weather station

Lake City. 3/24 No images on sd cards. Reset. Sven checked logger connections. Some debris on shelter roofs between cap and corrugated panels.

2023-01-13

We have not installed R2G5 shelter at Hancock for the last two seasons due to lack of stand from stressing crop. Todd/Chris attempted replanting in 2021.

2023-01-10

hancock 5/23/2022 8/29/2022 bcse 5/24/2022 8/30/2022 escanaba 5/24/2022 8/31/2022 rhinelander 5/24/2022 8/17/2022 lake-city 5/25/2022 8/18/2022 lux-arbor 5/27/2022 8/29/2022 2022 Shelter install and removal dates.

2022-09-15

Lake City, MC MLE, species composition transects - Alyssa, Emily, Helen, Gloria

2022-09-14

Escanaba, MN MLE, species composition transects - Alyssa, Emily, Helen, Gloria

2022-09-13

Rhinelander, WN MLE, species composition transects - Alyssa, Emily, Helen, Gloria

2022-09-12

Hancock, WC MLE, species composition transects - Alyssa, Emily, Helen, Gloria

2022-08-23

Removed shelters with Dave at Lux Arbor 11AM-1PM. R2-4 switchgrass quite stressed (for the first time I believe). Did not have RH/T sensors, supply chain issues this year. Removed and downloaded camera SD cards. Cards not replaced yet. G10 camera was on the ground.

2022-08-18

Lake City MLE drone flight for multispectral and visual (100\’ @ 80% side overlap). Also oblique images for catalog. Under MLE SD Cards file Kevin Docs. Removed shelters, RH/T sensors. Saved camera SD card data and returned (DW). Stress under R1 switchgrass.

2022-08-17

Escanaba stop. Kept shelters in, Todd to remove next week. Removed RH/T sensors from rep 1 and 3 inside outside shelters. Rain gauge okay. Camera cards downloaded and returned. Drone flight for multispec and visual 80\’, 80% side lap (same as others). Camera on G5 was on the ground. Repaired. Repaired gutter on G5R3.

2022-08-17

Rhinelander MLE stop. Removed all shelters. Reset cameras, checked asset tags, rain gage, removed RH/T sensors. R1 out did not respond. Drone flight for multispec and visual all. All data from cameras and drone saved on Kahmark computer\documents\MLE SD Cards folder. Shelters out by 11CDT.

2022-06-06

Shelters in on BCSE 5/24 3PM, Lux Arbor 5/27 11 AM, Hancock 5/23, Rhinelander 5/24, Escanaba 5/24, Lake City 5/25. Hancock, Escanaba, Rhinelander, and Lake City have RH/T sensors. Lux and BCSE no sensors available.

2022-05-26

Lux shelters installed. 12PM. No sensors installed (none available from Onset to replace broken units).

2022-05-21

Deployed Lake City RH/T sensors on G5 R1 and R4. Switched from R3 (previous) to R4 to see if stronger switchgrass stand and tree-line impact RH/T under shelter.

2022-05-20

Rain gauge at Lake City was clean. Sensor f on logger 125 was hogging the bus. I disconnected it. Collected images from the trail cams. Kevin deployed humidity sensors checked and repaired shelters (Kevin and Dave)

2022-05-20

Repaired rep 3 shelter. Sensors added in and out shelters.in R1 and E3. Posts in 0N. Attached hose, left screws. New battery on rep3 logger. Checked weather station and rain gauge all northern and central sites.

2022-05-19

RH/T sensors installed on Hancock G5 R3 shelter and out in 0N, and R4 shelter only. Did not silicone logger. Start logging at 5/19/22 14:15 CDT (central daylight savings time). No shelter install on R2 once again. R3 shelter area seems to have treatment effect, looking more sparse.

2022-05-19

checked raingauge at Escanaba. Funnel was clean.

2022-05-19

Humidity sensor started at Rhinelander at 18:30 CDT

2022-05-19

checked G8R3-10 profile probe. No response. Reseated all the cables. No change. Checked for visible breakage and found none.

2022-01-01

Replace weather station battery at Lux Arbor with a used battery from the field lab. Original battery was at 7.6V

2021-09-21

Lux 1, pakbus 113 shelter probe 2 is disconnected, but the note says that it works. I hooked it back up and all of the probes still responded.

2021-09-21

lux 2 pakbus 113. shelter sensors 1 and 2 were not working. Checked signal, power and ground connections all were tight and reconnecting did not bring the sensors back up. I don\’t have the handheld unit with me to check.

2021-09-20

MLE species composition and heights at Lux Arbor (MS) - Alyssa, Anna, Ellen

2021-09-17

MLE data sheet organization, identify plants - Stacey

2021-09-16

MLE species composition and heights at Lake City (MC) - Alyssa, Anna, Ellen, Stacey

2021-09-15

MLE species composition and heights at Escanaba (MN) - Alyssa, Anna, Ellen, Stacey

2021-09-14

MLE species composition and heights at Rhinelander (WN) - Alyssa, Anna, Ellen, Stacey

2021-09-14

Kevin removed shelters from the BCSE at 12:15 today.

2021-09-13

MLE species composition and heights at Hancock (WC) - Alyssa, Anna, Ellen, Stacey

2021-09-02

removed shelters from Lux Arbor around 3 pm.

2021-09-01

Raingauge at Rhinelander was clean. Repaired ground on 118 logger.

2021-08-31

Removed shelters from Lake City in the morning. Checked raingauges at Lake City and Escanaba, both were clean.

2021-06-22

Checked lux arbor mle rain gauge. All clear

2021-06-17

MLE trip from 6/14-6/16 Kevin and Sven. Installed BCSE and Lux shelters with Zaman on 6/11, Kevin adn Sven Lake City on 6/14, Escanaba and Rhinelander on 6/15, Hancock on 6/16. Installed RH/T sensors both under shelter and ambient 2 per at each site. Used tpost to install ambient in 0N approximately equivalent to shelter position and height. Now two ambient and two shelter sensors each site accept Lux and Hancock. Two sensors failed over last fall/winter. Changed solar panel angle, checked weather station, rain gage, and downloaded visual imagery from trail cams (cameras not keeping time well). Collected gravimetric samples from ambient vs shelter, a composite of four samples 0-10cm and 10-25. Only 0-10 at Hancock (too sandy). New sanitation station at Rhinelander!! Follow signage. Collected MMPRNT samples from all site (24 at each) LC, ESC, RHN, HAN.

2021-06-17

Lake City R1 shelter roof was damaged. The northwest panel was torn away. Repaired as best we could with half panels left at site. Dirty, but solid.

2021-06-16

hancock G5 camera is gone. replaced with a spare. Sampled G5-R3 around 9 am CDT. Sampled G5-R2 around 10 am. Radio communication stopped. I swapped the weather station radio with the G5R3 radio and it started working again.

2021-06-15

sampled gravimetric soil samples from G5 Escanaba around 10:00 and Rhinelander between 14:00 and 15:00 local time. Checked weather station. rain funnel clear at Rhinelander, plugged at escanaba.

2021-05-25

Kurt, Steve, Phil, Kevin, Sven decided to hold off installing all MLE shelters until week of June 14 based on current drought and late start of switchgrass growth due to cool weather.

2021-03-11

Lux Arbor MLE: All probes including shelter and profile that are connected to logger 113 were down. Shelter probes have been down before winter 2020. Sven and I checked these probes. Started checking first shelter probes by removing each probe (removing signal cable) and then connected to A200. All shelter probes were reading by A200. However, connecting probe id 2 back to the logger caused everything to fail again despite this probe being communicated with A200. This probe 2 is disconnected from the logger and labelled it separately with insulation tape. After this diagnosis. all profile probes are started to communicate as confirmed by Sven. There is one profile probe (id e) that is still missing that I disconnected on June 4, 2020. Refer log details of June 4, 2020 for probe id e. Collected images from cameras at Lux.

2020-11-04

Replaced a surface probe (id is \”g\”) in G5R3 in the logger 113 after surface probe was cut by harvesting machine at Lux. Replaced probe is working and there is data flow. Sprayed chemical to kill wasp hive that formed inside the logger box near the entrance port.

2020-09-27

All tested probes of G8R2 and G10 R2 (logger 115) at Lux MLE are backed and reporting values.

2020-09-25

With Sven help, probes of G10R2 and G8R2 (logger 115) at Lux Arbor were troubleshooted. All probes were first tested by A200 interface and later by remote logging. Probes found communicating by A200. By remote logging, three probes were first tested (by disconnecting other 3 probes from logger), and later tested other probes by connecting each one at a time. By remote logging, all probes found communicating. Yet, only nan’s are reporting. More updates will follow.

2020-09-21

Checked G8R2 and G10R2 plots in the logger 115 at Lux Arbor MLE. This logger was down since 28 Aug, there were not even Nan\’s since then. Before anything, solar controller, logger and radio were checked and they were powered. Logger rebooted and again reconnected (remotely by Sven), seems all probes are sending Nan\’s. However, probes f and d were tested randomly at the field by A200, all reading SWC. More details will follow. Correction: The logger was working and the probes are sending NaN\’s. But the 3 month dataset in weather/tables excludes any values with NaN\’s to avoid having to deal with NaN\’s in the shiny app. –sven

2020-09-14

Kevin reports that the raingauge at Lake City was pugged and the funnel was full of water.

2020-09-02

Collected Escanaba RH-T sensor data on 8/31, Rhinelander 9/1, and Hancock on 9/1. At Hancock we had only one ambient sensor and that sensor was full of water upon collection. No data so need to use weather station as a substitute. Will compile all data into long form and submit to Sven.

2020-08-21

Downloaded RH/Temp data from sensors at Escanaba, Rhinelander and Hancock over the last few days. The only ambient sensor at Hancock was full of water so we will have no direct comparison of under shelter to ambient. Will combine data into a long format and submit to Sven. Collect Lux and Lake City next week.

2020-07-08

rain gauge funnel was clean at Escanaba, no spiders either. Installed rebooting device. Shelter on G5R3 was blown out of position. The north west corner was almost where the north east corner was. I moved it back into postion

2020-07-07

checked rain gauge funnel at rhinelander, it was clean

2020-07-07

cleaned screen on raingauge at Hancock. Funnel was clean. Installed reboot device.

2020-06-30

reconnected radio on 113 logger at Lux Arbor. Starting to see some drought effect in some of the replicates.

2020-06-29

fixed firmware issue at Lake City. Some drought stress under the shelters. Rain gauge funnel was clean

2020-06-24

Tried to install the surface probe (0-30 cm) in G5R4 which is connected to logger 113 (located in G5R3). While removing broken sensor (probe id \”e\”), I realized this probe has got a cable length of 150 feet and run through the ditch along with other cables (distance from G5R3 plot to G5R4- roughly 100-125 feet). This probe can be installed either by removing broken sensor cable from the ditch or install in a newer ditch. Kevin suggested adding this sensor either way in fall time when no canopy closure.Therefore, there is no soil water data from 0-30 cm depth in G5R4.

2020-06-04

Kevin and I, checked the profile probes in the logger 113 at Lux Arbor MLE. Kevin connected the logger program directly to the logger, then disconnected all profile probes (only green signal cable) from the logger and reconnected them one by one and checked. All profile probes connected to logger are good, except probe id \”e\” failed to communicate. The probe id \”e\” is 0-30 cm probe from G5R4 plot. I will soon replace this failed probe. It may be likely that thunderstorm hit this probe, and caused other profile probes from not working.

2020-06-02

For SDI12 troubleshooting: Open Loggernet, setup, make sure COM port correct, Use male/female RS232 to USb adapter, select proper logger and comport. Go to connect, and connect. Select datalogger from top bar and terminal emulator. Click on open terminal. enter twice. Should read CR800>. Then type SDI12 exactly. return. should show available control ports. Choose 1 or 2 then type in address and check for response a!

2020-05-30

Todd Martin installed Escanaba shelters.

2020-05-29

Todd martin installed Hancock and Rhinelander shleters on MLE G5 plots.

2020-05-28

Checked the logger 113 that connects the profile probes of g8r1, g10r4, g5r4 and g5r3. These probes were stopped working at same time on 14 May 2020 \@3:00. Thunderstorm were reported on 13th and 14th May 2020 around Kalamazoo. Shelter probes are working. Checked the power supply from voltage controller to terminal blocks (that has connected with probes) and from terminal blocks to logger ports (Shelter probes connected to C1, profile probes connected to C3), power looks OK. Also checked the grounding that also look OK. Sven said, this logger reported \”nan\” for profile probes. Kevin and Sven suggested to check each probes (using A200) by disconnecting from logger and then reconnecting to it. Well, the day rained a lot. We will get this work tomorrow when sunny. Downloaded the images from Camera. Since the setting in those camera\’s were not right, so there were photos but with different date and time stamps which is not right. Kevin set all the camera parameters. This camera needs alkaline batteries. Camera facing the poplar plot (G8R3) that was tied to logger box at G5R2, is moved and tied to deer fence post of G8R3. It is located on the north-west side of the fence.

2020-05-26

Checked 113 logger at Lux Arbor MLE site. Profile soil water probes connected to this logger has been stopped working since last week storm. However, shelter probes are continue to work. Profile probes in this logger that were hit by storm includes all depths in G8R1, G5R3, G10R4 and G5R4 plots. I checked the voltage and power supply from solar to logger and from terminal blocks to logger ports. There is a good power supply and cable connectivity. All shelter probes in the logger are working. I did not disconnect the profile probes nor the profile probes were checked by A200 device utility Interface. I feel, checking each probes with A200 interface is not really important at this time. I just rebooted the logger. If the probes does not return after rebooting, then I must come back and check the probes by A200.